# Forecasting Local Ionospheric Parameters Using Transformers

### v1

@@ Abstract
We present a novel method for forecasting key ionospheric parameters using transformer-based neural networks. The model provides accurate forecasts and uncertainty quantification of the F2-layer peak plasma frequency (foF2), the F2-layer peak density height (hmF2), and total electron content (TEC) for a given geographic location. It supports a number of exogenous variables, including F10.7cm solar flux and disturbance storm time (Dst). We demonstrate how transformers can be trained in a data assimilation-like fashion that use these exogenous variables along with naïve predictions from climatology to generate 24-hour forecasts with nonparametric uncertainty bounds. We call this method the Local Ionospheric Forecast Transformer (LIFT). We demonstrate that the trained model can generalize to new geographic locations and time periods not seen during training, and we compare its performance to that of the International Reference Ionosphere (IRI).
@@ 1 Introduction
Accurate and efficient modeling of Earth’s ionosphere has a significant impact on research and operational communities due to its effects on radio communications, radar performance, [1, 2, 3] and satellite drag [4]. Success in forecasting key parameters such as the F2 layer critical frequency (foF2) and height (hmF2) and the total electron content (TEC) allows one to anticipate and mitigate the impacts of ionospheric variability on such systems. Over the past decades, many modeling approaches have been developed to predict these ionospheric parameters with increasing accuracy and skill. These models may be broadly categorized as empirical, physics-based, and, more recently, machine learning methods.
@@
Empirical models often rely on extensive historical datasets to establish statistical relationships between ionospheric parameters and geophysical variables. The International Reference Ionosphere (IRI) model [5] is a widely used standard that provides monthly averages of various ionospheric parameters based on many decades of past observations. IRI has seen continual development and improvements over the years, adding a host of submodels used to capture specific aspects of the ionosphere such as the CCIR [6, 7] and URSI [8] foF2 models for representing the diurnal variations of the peak plasma density across the globe, the AMTB [9] and SHU-2015 [10] models for even more harmonic expansions of hmF2, and NeQuick 2 [11] for improved topside electron density accuracy and thus better estimates of TEC [12, 13]. So, while large empirical models like IRI continue to improve, the number of these available options needed to address each domain and source of variance in the ionosphere also grows, and choosing the appropriate settings may be prohibitive without expert knowledge of each submodel.
@@
The Physics-based models instead aim to incorporate the fundamental processes governing the dynamics of the ionosphere and its drivers by solving the highly complex sets of equations related to the plasma magnetohydrodynamics, electrodynamics, and chemistry. These often seek to simulate the ionospheric response to solar and geomagnetic inputs. The Thermosphere-Ionosphere-Mesosphere-Electrodynamics General Circulation Model (TIME-GCM) [14], the Coupled Thermosphere Ionosphere Plasmasphere Electrodynamics (CTIPe) model [15], and SAMI3 [16] are examples of such models, and while they offer detailed insights through highly accurate nowcast and hindcast, their ability to forecast is often limited, and forecast skill can drop precipitously over relatively short time windows, even with careful treatment of driving parameters such as solar and geomagnetic activity. Additionally, the simulations themselves can be computationally expensive, requiring high-performance computing to generate high-resolution or large numbers of simulations.
@@
Both empirical and physics-based methods require special treatment of drivers and parameterizations in order to maintain accuracy over a forecast period. Although there are efforts to develop better indices for these drivers that can be included in models [17], the impact on forecast skill is still largely determined by how well these drivers can inform future values of the parameters and not just how correlated they are in the past. This challenge is further complicated by the fact that these drivers, which often take the form of global indices, may interact with the forecasted parameters with unknown or varying time delays [18, 19, 20, 21, 22]. Thus, determining the complex interactions between multivariate time series of drivers and observations motivates the consideration of machine learning as a component in the next generation of forecast models.
@@
Learned latent representations can greatly simplify a time series through linearizations of the latent representation [23] or time delay embeddings [24]. However, the burden of choosing the correct latent dimension or time delay can make these methods cumbersome when applied to real-world datasets, and there is no straightforward means of validating that a given embedding will continue to work for future time series or, in the case of ionospheric forecasting, for new geographic locations. Nevertheless, recent advances in transformer-based neural networks have produced promising results [25, 26].
The parameters we wish to forecast are readily available through sounding measurements and subsequent inverse processing, yet the multiscale and noisy nature of these data makes them difficult to predict using conventional spectral decompositions [27]. We find that transformers allow us to overcome all of these challenges at once. Namely, we analyze the use of transformer networks and demonstrate their ability to ingest a complex mix of inputs, including geomagnetic and solar activity, to produce a probabilistic forecast of key ionospheric measures. Moreover, it is simple to also feed the transformer a naive prediction from an existing model such as IRI as well. Naive in this context, implying that the settings and parameters for IRI are not tuned by the user. Thus, the transformer learns to dynamically assimilate those data into the naive prediction. Another major drawback of current empirical and physics-based models is the lack of or difficulty in generating uncertainty quantification. Our use of transformers also allows us to take steps towards addressing this by way of multi-quantile regression. So, instead of the model producing point forecasts over the prediction horizon, we produce a set of quantiles that are learned from the data that capture empirical distributions of uncertainty in each target parameter. We call this model the Local Ionospheric Forecast Transformer (LIFT).
@@
The following sections will describe the data used to train this model, Section 2, the model architecture itself in Section 3, the results of these forecasts, Section 4, and conclusions for future directions, Section 5. The results of this model suggest that transformers may play an essential role in the next generation of data-driven ionospheric models, and while they lack the interpretability of the full-physics systems, their ability to transform new data into forecasts makes them a power tool for operational communities.
@@ 2 Data
Our model is trained using data from several openly available sources. The observational data for the target parameters, foF2, hmF2, and TEC are obtained from the Lowell GIRO Data Center digital ionogram database (DIDBase) [28]. These parameters are determined from raw sounder data by way of an autoscaling and inversion algorithm, ARTIST5 [29]. The sounder data acquisition and subsequent parameter computation has a number of potential sources of error, which ARTIST5 attempts to quantify using a confidence score; we filter out any poorly inverted parameters using a heuristically chosen threshold, retaining only values with confidence $\geq 70$. Higher confidence scores do not necessarily correspond to better autoscalings. However, this threshold has been assessed to cover autoscalings that are accurate enough for the foF2 and hmF2 parameters; see [30] for a more detailed analysis of the confidence scores of ARTIST5.
@@
Often, forecast models are tested by evaluating them on a held out data set consisting of measurements from future time steps that the model has not seen. Unfortunately, for a localized ionospheric forecast, this tells us little about how the model will perform data measured at a new location. Instead, we elect to perform our testing on data from the same time periods as the training data, but from completely new geographic regions, i.e., we hold out entire stations’ data for testing. Figure 1 indicates which station locations were used for training and validation versus test evaluation in our experiments. Although most available sounding measurements are located in the Northern hemisphere, a few test stations are located near the equatorial region and the lower hemisphere; we reserve these held out stations entirely for testing, which allows us to demonstrate our model’s ability to generalize to novel geographic locations, and new parts of the ionosphere, after training. This makes our LIFT model more useful in practice, since a user is not required to retrain or tune the model in order to forecast at a new station. Because we are training a local forecaster, all locations are kept in geographic latitude and longitude, and no conversions to magnetic coordinates were needed.
Our model ingests windows of data that are 72 hours in length, and generates a forecast for the subsequent 24 hours; these must be contiguous periods of observations, since we wish to learn sequential temporal relationships between past measurements and covariates and future observations. After removing data with poor ARTIST5 confidence autoscalings, we further downselect to only contiguous time series from each station that contain more than the minimum number of points needed by the model, i.e., 96 hours. Each of these contiguous time series is given a unique segment identifier and this process is performed separately for the training, validation, and test sets. Figure 3 provides a summary of the number of unique sounder observations over time for the training, validation, and test sets. The total number of points and contiguous segments are reported in Table 1.0
@@
missingNumber of points missingNumber of 96-hour segments
Training 9,336,446 528,413
Validation 642,821 38,494
Testing 1,481,978 76,220
@@
Each data segment is then paired with a corresponding set of drivers that are used as exogenous inputs to the model. These drivers include sunspot number (SSN), the Kp index (Kp), and F10.7, which were obtained from the GFZ German Research Center for Geosciences [31], and the Dst index from the WDC for Geomagnetism, Kyoto [32]. Together, these indices represent a set of additional covariates that can be measured or obtained alongside new sounding data, but whose interactions with the ionospheric parameters may be too complex, nonlinear, or lagged to be expressed as closed analytic expressions. Figure 4 illustrates the affiliated time series of these indices for the Boulder, CO, USA ionosonde station.
@@ 3 Methods
where $c$ and $p$ are the number of context and prediction points, respectively. The matrix $\mathbf{X}$ contains the foF2, hmF2, and TEC parameters observed from an ionosonde, thus $d_{1}=3$, $\mathbf{Y}$ corresponds to the data in Figure 4 and consists of the exogenous or forcing variables that we expect to influence the forecast, so $d_{2}=4$, and $\mathbf{Z}$ contains data for computable covariates, including a time vector and the naïve IRI prediction, which we describe in more detail in Sections 3.3 and 3.4, resulting in $d_{3}=13$.
@@
The infimum in this definition is necessary only if the CDF happens to be constant on some interval, in which case $Q_{x_{t+k;j}\mid\mathbf{X}_{t-c:t},\mathbf{Y}_{t-c:t},\mathbf{Z}_{t-c:t+p}}(\tau)$ would not be well defined [33, 34, 35].
@@
The linear component is used to predict the simple linear trends that we expect to see in each of the target parameters by generating predictions for the next $p$ time steps based solely on the previous $c$ time steps. Furthermore, it predicts all $p$ time steps simultaneously, similar to sequence-to-sequence models [36], and so we refer to it as a direct multistep linear model. We fit independent linear models for each of the $j$ target variables (foF2, hmF2, TEC),
@@
with the superscript $L$ denoting this prediction comes from the linear model. The weight matrices $\mathbf{W}_{j}\in\mathbb{R}^{p\times(c+1)}$ and bias vectors $\mathbf{b}_{j}\in\mathbb{R}^{p}$ are fit using the Adam optimizer [37] alongside the transformer network. So, this model only generates a linear approximation for a vector of future values from a vector of past values. This forecast will have residual errors,
@@
Conceptually, the transformer acts as a nonlinear correction to the simple linear model that incorporates all available data, including drivers and naïve IRI-based climatology, but without the need to specify explicit functional relationships to the target parameters foF2, hmF2, and TEC. Figure 5 provides a concrete example of the output of each of the LIFT model components and Figure 6 illustrates the complete end-to-end architecture of our LIFT model architecture.
@@
Figure 6: Complete LIFT model diagram. All components of the model related to the transformer are shaded green, while the linear model is indicated by red, the inputs in yellow, and the outputs in blue. The stacks each correspond to an individual forecasted parameter (foF2, hmF2, and TEC), so there are three sets of outputs from the linear models and the transformer. The final forecast data at the top is a matrix containing all aggregated outputs.

### v2

# arXiv:2502.15093v2[physics.space-ph] 27 Feb 2026

@@ Abstract
We present a novel method for forecasting key ionospheric parameters using transformer-based neural networks. The model provides accurate forecasts and uncertainty quantification of the F2-layer peak plasma frequency (foF2), the F2-layer peak density height (hmF2), and total electron content (TEC) for a given geographic location. It includes a number of exogenous variables, including F10.7cm solar flux and disturbance storm time (Dst). We demonstrate how transformers can be trained in a data assimilation-like fashion that uses these exogenous variables along with naïve predictions from climatology to generate 24-hour forecasts with nonparametric uncertainty bounds. We call this method the Local Ionospheric Forecast Transformer (LIFT). We demonstrate that the trained model can generalize to new geographic locations and time periods not seen during training, and we compare its performance to that of the International Reference Ionosphere (IRI) using CCIR coefficients.
@@ 1 Introduction
Accurate and efficient modeling of Earth’s ionosphere has a significant impact on research and operational communities due to its effects on radio communications, radar performance (?, ?, ?, ?), and satellite drag (?, ?). Success in forecasting key parameters such as the F2-layer critical frequency (foF2) and height (hmF2) and the total electron content (TEC) allows one to anticipate and mitigate the impacts of ionospheric variability on such systems. Over the past decades, many modeling approaches have been developed to predict these ionospheric parameters with increasing accuracy and skill. These models may be broadly categorized as empirical, physics-based, and, more recently, machine learning methods.
@@
Empirical models often rely on extensive historical datasets to establish statistical relationships between ionospheric parameters and geophysical variables. The International Reference Ionosphere (IRI) model (?, ?) is a widely used standard that provides monthly averages of various ionospheric parameters based on many decades of past observations. IRI has seen continual development and improvements over the years, adding a host of submodels used to capture specific aspects of the ionosphere such as the CCIR (?, ?, ?) and URSI (?, ?) foF2 models for representing the diurnal variations of the peak plasma density across the globe, the AMTB (?, ?) and SHU-2015 (?, ?) models for harmonic expansions of hmF2, and NeQuick 2 (?, ?) for improved topside electron density accuracy and thus better estimates of TEC (?, ?, ?). While large empirical models like IRI continue to improve, the number of available options needed to address each domain and source of variance in the ionosphere also grows, and choosing the appropriate settings may be prohibitive without expert knowledge of each submodel.
@@
Physics-based models instead aim to incorporate the fundamental processes governing the dynamics of the ionosphere and its drivers by solving highly complex sets of equations related to plasma magnetohydrodynamics, electrodynamics, and chemistry. These models seek to simulate the ionospheric response to solar and geomagnetic inputs. The Thermosphere-Ionosphere-Mesosphere-Electrodynamics General Circulation Model (TIME-GCM) (?, ?), the Coupled Thermosphere Ionosphere Plasmasphere Electrodynamics (CTIPe) model (?, ?), and SAMI3 (?, ?) are examples of such models. While they offer detailed insights through highly accurate nowcast and hindcast, their forecast skill can drop precipitously over relatively short time windows, even with careful treatment of driving parameters such as solar and geomagnetic activity. Additionally, the simulations themselves can be computationally expensive, requiring high-performance computing to generate high-resolution or large numbers of simulations.
@@
Both empirical and physics-based methods require special treatment of drivers and parameterizations in order to maintain accuracy over a forecast period. Although there are efforts to develop better indices for these drivers (?, ?), the impact on forecast skill is largely determined by how well these drivers can inform future values of the parameters and not just how correlated they are in the past. This challenge is further complicated by the fact that these drivers, which often take the form of global indices, may interact with the forecasted parameters with unknown or varying time delays (?, ?, ?, ?, ?, ?). A further limitation of current empirical and physics-based models is their inability to provide robust uncertainty quantification. The dual challenge of capturing complex, time-delayed driver interactions and providing reliable uncertainty estimates motivates the integration of machine learning into next-generation forecasting systems.
@@
Many machine learning models operate on learned latent representations of the data. These learned spaces can greatly simplify a time series through linearizations of the latent variable dynamics (?, ?) or time delay embeddings (?, ?). However, the burden of choosing the correct latent dimension or time delay can make these methods cumbersome when applied to real-world datasets, and there is no straightforward means of validating that a given embedding will continue to work for future time series or for new geographic locations. Recent advances in transformer-based neural networks offer a promising alternative (?, ?, ?, ?, ?, ?).
The use of transformers for forecasting in space-weather applications has grown significantly in recent years. Current work primarily uses recurrent networks and attention-based architectures for ionospheric forecasting, with the bulk of activity focused on global TEC maps. For instance, sequence-to-sequence transformers trained on global ionospheric maps (GIMs) from the Center for Orbit Determination in Europe (CODE) achieve stable multi-step forecasts up to 48 hours, with root-mean-squared error (RMSE) reported near 1.8 TECU (?, ?). Extensions that pair transformers with representation learning or denoising have also been developed. For example, an “improved transformer” that integrates deep denoising autoencoders to compress and clean GIMs reduces errors relative to conventional baselines and demonstrates better robustness during disturbed conditions (?, ?). Other efforts exploit autocorrelation mechanisms and synthetic data augmentation to capture long-range dependencies, as seen in a synthesis-style autocorrelation transformer that provides one-day GIM forecasts and outperforms operational one-day products on a 2018 benchmark (?, ?). Capturing spatial couplings explicitly has also proved beneficial. A graph-enabled spatiotemporal transformer (GEST) models station connectivity with graphical-model layers, while a time-transformer attention block handles temporal dependencies; this approach yields higher accuracy than prior deep models for multistep TEC prediction and improves storm-time performance (?, ?). More hybrid designs have appeared, combining convolutional encoders for local spatial patterns with transformer temporal modules, such as convolutional attentional image time-sequence transformers for TEC maps (?, ?). More recently, autoformer variants with decomposition layers have been applied to global TEC forecasting (?, ?), and deep learning methods including LSTM and GRU have been developed toward operational forecasting services with incremental learning capabilities (?, ?). Related spatiotemporal encoder–decoder variants claim state-of-the-art results across solar-cycle phases.
These studies collectively point to several design levers that matter for TEC forecasting, including explicit spatial modeling (graphs or CNNs), long receptive fields via attention, and incorporation of drivers such as F10.7, Kp, Dst, and solar-wind parameters. For foF2, recent transformer applications adopt long-sequence variants tailored to single-station forecasting. Informer architectures (?, ?), which use probabilistic sparse attention and distillation to reduce complexity, improve 5- to 48-hour foF2 forecasts over recurrent (LSTM) and empirical (IRI) models at both middle and low latitudes (?, ?). Reported implementations train on multi-year ionosonde records with exogenous inputs (F10.7, Dst, Kp), maintain skill during geomagnetic storms, and reduce RMSE across the forecast horizon compared with tuned recurrent baselines (?, ?). These results indicate that attention mechanisms help recover diurnal and seasonal structure while adapting to storm-time deviations when given driver indices. By contrast, explicit transformer formulations for hmF2 remain sparse in the open literature; most studies targeting this parameter still rely on feedforward neural networks or recurrent models and report improvements over IRI or simple autoregressive baselines (?, ?).
However, existing transformer models often use multiple attention layers with large embedding dimensions, which drastically increases model size relative to the available data. For instance, the Informer foF2 forecast model in ? (?) uses 10 encoder layers and 8 decoder layers, each with embedding dimensions of 512 and feed-forward networks with a 2048 inner dimension, resulting in tens of millions of trainable parameters. This is especially relevant for spatiotemporal data such as GIMs or ionosonde time series, as batching generates highly correlated samples. Consequently, there is a gap in the existing literature for parsimonious transformer-based forecasters that jointly predict foF2, hmF2, and TEC, leverage exogenous drivers, and provide calibrated uncertainty; cf. ? (?) for a comprehensive review demonstrating this gap.
We address this gap with the Local Ionospheric Forecast Transformer (LIFT). Our approach differs from prior work in several respects. First, we use a compact, single-layer attention network combined with a simple linear baseline. The linear component captures a large fraction of the variation in ionospheric parameters for short horizons (e.g., 24 hours), while the transformer provides data-adaptive corrections and multi-quantile outputs for uncertainty quantification. With a single layer in both the encoder and decoder and an embedding dimension of 128, the model contains only roughly 370K parameters. Second, the model is localized, i.e. trained on single-station inputs rather than GIMs, yet can generalize to new geographic locations. This is an important use case in edge-computing environments when the only available data may come from a local ionosonde, but accurate regional forecasting is critical. Training on single-station inputs also allows the model to leverage multiple decades of dense (15-minute cadence) ionosonde observations with lower computational overhead. Third, it is straightforward to incorporate naïve predictions from an existing climatological model as an additional transformer input, providing a baseline that can be adaptively corrected.
Our training procedure also differs from prior work. Many existing studies evaluate models by holding out the final one to two years of data for testing and training on earlier data. While this is common in time-series forecasting, the periodicity of the solar cycle and its influence on ionospheric dynamics mean such a test set evaluates performance only for a particular phase of solar activity. We generate training and test splits via a striping method (?, ?, ?), using 23 years of data that span roughly two full solar cycles to produce samples throughout the entire period. The test set uses geographic locations held out entirely from training, demonstrating that the model can generalize to new locations without the need for complex graphical models.
To benchmark our forecasts, we compare LIFT against PyIRI (?, ?), a Python implementation of the IRI model optimized for computational efficiency. As described in ? (?), earlier versions of PyIRI (e.g., v0.0.4) incorporated a mixture of traditional IRI and NeQuick components using NeQuick-derived formulations for parameters such as bottomside and topside thickness, topside shape factors, and other profile characteristics. In this study, we use PyIRI v0.1.5, which has been updated to closely follow the standard IRI formulation and no longer relies on NeQuick elements. For the hmF2 parameter, PyIRI provides several options, including SHU2015, AMTB2013, and BSE-1979 (?, ?); we use the SHU2015 option in this work. A detailed comparison of these submodels is available in ? (?). PyIRI’s computational speed enables the large-scale simulations required to generate benchmark forecasts across our training and test datasets. As the most widely used empirical ionospheric model, IRI serves as a trusted standard against which new forecasting methods are commonly evaluated.
@@
The following sections present the data used to train the model (Section 2), the model architecture (Section 3), the forecasting results (Section 4), and conclusions with directions for future work (Section 5). The results indicate that transformers may become a central component of next-generation data-driven ionospheric models. Although they lack the interpretability of full-physics systems, their ability to convert new data into accurate forecasts makes them a powerful tool for operational applications.
@@ 2 Data
Our model is trained using data from several openly available sources. The observational data for the target parameters, foF2, hmF2, and TEC are obtained from the Lowell GIRO Data Center digital ionogram database (DIDBase) (?, ?). These parameters are determined from raw sounder data by way of an autoscaling and inversion algorithm, ARTIST5 (?, ?). The sounder data acquisition and subsequent parameter computation have a number of potential sources of error, which ARTIST5 attempts to quantify using a confidence score; we filter out any poorly inverted parameters using a heuristically chosen threshold, retaining only values with confidence $\geq 50$ and excluding 55. Higher confidence scores do not necessarily correspond to better autoscalings for all parameters. However, this threshold has been assessed to cover autoscalings that are accurate enough for the foF2 and hmF2 parameters; see (?, ?) for a more detailed analysis of the confidence scores of ARTIST. Figure 1 shows the distribution of confidence scores over the entire training, validation, and test sets. A future effort to train a model such as LIFT on pristine F2 and TEC parameters would be of great interest. Unfortunately, this quality of data is generally only available via manually scaled ionograms, which are not readily available in the open literature and is evident in Figure 1 by the relatively few number of code 999 observations indicating manually scaled profiles.
Figure 1: Distributions of the ARTIST confidence scores across the training, validation, and test sets.
@@
Often, forecast models are tested by evaluating them on a held-out data set consisting of measurements from future time steps that the model has not seen. Unfortunately, for a localized ionospheric forecast, this tells us little about how the model will perform on data measured at a new location. Instead, we test on data from the same time periods as the training data but from completely new geographic regions; i.e., we hold out entire stations’ data for testing. Figure 2 indicates which station locations were used for training and validation versus test evaluation in our experiments. Although most available sounding measurements are located in the Northern Hemisphere, a few test stations are located near the equatorial region and the Southern Hemisphere; we reserve these held-out stations entirely for testing. This experimental design allows us to evaluate the model’s ability to generalize to novel geographic locations after training. As quantified in Table 4, the LIFT model produces skillful forecasts across low, mid, and high geomagnetic latitude bands, though performance varies with ionospheric regime. This approach eliminates the need for retraining or tuning when forecasting at a new station, making the LIFT model practical for operational deployment at locations not represented in the training data. Because we are training a local forecaster, all locations are kept in geographic latitude and longitude, and no conversions to magnetic coordinates were needed.
Our model ingests windows of data that are 72 hours in length and generates a forecast for the subsequent 24 hours; these must be contiguous periods of observations, since we wish to learn sequential temporal relationships between past measurements and covariates and future observations. After removing data with poor ARTIST5 confidence autoscalings, we further downselect to only contiguous time series from each station that contain more than the minimum number of points needed by the model, i.e., 96 hours. During this downselection process, a 96-hour window is moved across each variable of the entire dataset. If a 96-hour segment has only a few observations missing (<10<10), then we use 1-D piecewise polynomial interpolation to fill in those gaps. We ensure these missing data points do not all lie on the endpoints. Any 96-hour segment that has >10>10 missing points in either the foF2, hmF2, or TEC parameter from the sounder is not considered, and the window is moved forward.
Each of these contiguous time series is given a unique segment identifier, and this process is performed separately for the training, validation, and test sets. Figure 4 provides a summary of the number of unique sounder observations over time for the training, validation, and test sets. The total number of points and contiguous segments are reported in Table 1.
@@
Number of points Number of 96-hour segments
Training 9,383,186 530,677
Validation 585,725 33,697
Testing 1,494,598 76,894
@@
Each data segment is then paired with a corresponding set of drivers that are used as exogenous inputs to the model. These drivers include sunspot number (SSN), the Kp index (Kp), and F10.7 cm solar flux, which were obtained from the GFZ German Research Center for Geosciences (?, ?), and the Dst index from the WDC for Geomagnetism, Kyoto (?, ?). Together, these indices represent a set of additional covariates that can be measured or obtained alongside new sounding data, but whose interactions with the ionospheric parameters may be too complex, nonlinear, or lagged to be expressed as closed analytic expressions. Figure 5 illustrates the affiliated time series of these indices for the Boulder, CO, USA ionosonde station.
@@
All input variables were aligned to a common 15-minute temporal cadence. Ionosonde observations of foF2, hmF2, and TEC were obtained at this native sampling rate from the GIRO repository. Geomagnetic and solar indices that are reported at coarser temporal resolutions were forward-filled to match this cadence: Dst values (1-hour resolution) were repeated across four consecutive time steps, Kp values (3-hour resolution) across twelve time steps, and SSN (daily resolution) across all 96 time steps in each day. This forward-filling approach is consistent with operational practice, where the most recently available index value is applied until a new measurement is reported. Covariates that can be computed analytically, including solar zenith angle and PyIRI climatological predictions, were evaluated directly at each 15-minute time step. The transformer’s self-attention mechanism learns to weight inputs according to their relevance to the forecast, so this temporal alignment does not require uniform information content across all time steps; the model can learn to attend preferentially to time steps at which coarse-cadence indices are updated.
@@ 3 Methods
where $c$ and $p$ are the number of context and prediction points, respectively. The matrix $\mathbf{X}$ contains the foF2, hmF2, and TEC parameters observed from an ionosonde, thus $d_{1}=3$, $\mathbf{Y}$ corresponds to the data in Figure 5 and consists of the exogenous or forcing variables that we expect to influence the forecast, so $d_{2}=4$, and $\mathbf{Z}$ contains data for computable covariates, including a time vector and the naïve PyIRI prediction, which we describe in more detail in Sections 3.3 and 3.4, resulting in $d_{3}=13$.
@@
The infimum in this definition is necessary only if the CDF happens to be constant on some interval, in which case $Q_{x_{t+k;j}\mid\mathbf{X}_{t-c:t},\mathbf{Y}_{t-c:t},\mathbf{Z}_{t-c:t+p}}(\tau)$ would not be well defined (?, ?, ?, ?).
@@
The linear component is used to predict the simple linear trends that we expect to see in each of the target parameters by generating predictions for the next $p$ time steps based solely on the previous $c$ time steps. Furthermore, it predicts all $p$ time steps simultaneously, similar to sequence-to-sequence models (?, ?), and so we refer to it as a direct multistep linear model. We fit independent linear models for each of the $j$ target variables (foF2, hmF2, TEC),
@@
with the superscript $L$ denoting this prediction comes from the linear model. The weight matrices $\mathbf{W}_{j}\in\mathbb{R}^{p\times(c+1)}$ and bias vectors $\mathbf{b}_{j}\in\mathbb{R}^{p}$ are fit using the Adam optimizer (?, ?) alongside the transformer network. So, this model only generates a linear approximation for a vector of future values from a vector of past values. This forecast will have residual errors,
@@
Conceptually, the transformer acts as a nonlinear correction to the simple linear model that incorporates all available data, including drivers and naïve IRI-based climatology, but without the need to specify explicit functional relationships to the target parameters foF2, hmF2, and TEC. Figure 6 provides a concrete example of the output of each of the LIFT model components, and Figure 7 illustrates the complete end-to-end architecture of our LIFT model architecture.
@@
Figure 7: Complete LIFT model diagram. All components of the model related to the transformer are shaded green, while the linear model is indicated by red, the inputs in yellow, and the outputs in blue. The stacks each correspond to an individual forecasted parameter (foF2, hmF2, and TEC), so there are three sets of outputs from the linear models and the transformer. The final forecast data at the top is a matrix containing all aggregated outputs. We include the $N$ and $M$ blocks for the transformer blocks for completeness. The LIFT model uses $N=M=1$.